OpenStudy (anonymous):

Is the following situation a binomial setting? explain how it does or does not meet the conditions of a binomial setting. Draw a single card from a standard deck of cards. Observe the card and then place it. Count the number of times you draw a card like this until you get a seven

OpenStudy (turingtest):

what is the probability of drawing a seven each time?

OpenStudy (anonymous):

1/52

OpenStudy (turingtest):

actually there are four 7's in a standard deck, so 4/52 what is the probability of not getting seven?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

0.077

OpenStudy (turingtest):

what is the probability of not getting seven?

OpenStudy (anonymous):

is it 1/52 or 1/48

OpenStudy (turingtest):

it is neither. the probability of something happening plus the probability of it not happening must always be equal to 1.

OpenStudy (anonymous):

only two choices, success= seven failure= any other card each draw is independent because the card is replace each time. Draw until you observe a 7 the probability of success, p=4/52=1/3

OpenStudy (turingtest):

right, so does that fit the definition of a binomial probability?

OpenStudy (anonymous):

it is geometric

OpenStudy (anonymous):

BINOMIAL DISTRIBUTION, POPULATION PROPORTION n = NUMBER OF TRIALS [ 7 ] (sample size) k = NUMBER OF SUCCESSES [ 1 ] (from 0 up to and including k NUMBER OF SUCCESSES) p = POPULATION PROPORTION [ 0.08 %]

OpenStudy (turingtest):

yes it is geometric but that wasn't part of the question so I ignored it. where do you get 20 trials from?

OpenStudy (turingtest):

where do you get 7 trials from?

OpenStudy (anonymous):

does or does not meet the conditions of a bionomial. until you get a seven

OpenStudy (anonymous):

Bionomial settingis failure or success

OpenStudy (turingtest):

what is the definition of a binomial probability distribution?

OpenStudy (turingtest):

right, and do we have that on each trial or not?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

on each trial, do we not have a chance of success p=4/52 and failure=1-p ?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

is that binomial?

OpenStudy (anonymous):

so it can be bionomial because it met one of the four conditions

OpenStudy (turingtest):

Again, you have to be careful about the definition of a binomial distribution. Binomial distributions have 2 parameters; the probability and the number of trials. Although each trial here is binomial, because we are drawing cards *until* the first 7 is found, the number of trials is not specified. Hence geometric probabilities are never the same as binomial ones.

OpenStudy (anonymous):

ok thanks